# Backstepping for Partial Differential Equations

### v2

@@ 3.4 Interconnected systems with a chain structure
The framework proposed by the representation (51) is comprehensive and can describe various interconnected systems. As shown in [21], using a technique referred to as folding (see [103, 20, 25] for details), it can describe complex networks composed of interconnected hyperbolic/ODEs subsystems. However, such a representation shadows the cascade structure between the different subsystems, which may lead to strong requirements when designing stabilizing controllers. In particular, the contributions presented in Section 3.3.3 avoid configurations where the PDE system is under-actuated or under-measured. Consequently, they fail to design stabilizing controllers for interconnected PDE systems when only one of the PDE is actuated. Consider the following interconnections of PDE of heterodirectional systems
@@ 3.5 Output regulation and disturbance rejection
A natural extension of the analysis of coupled PDE-ODE systems is the consideration of output tracking and disturbance rejection problems with a disturbance or reference model. In this case, when the disturbance and output trajectories can be modeled by a finite-dimensional exosystem, backstepping designs exist for the disturbance rejection and output tracking problems. The main technical difference with respect to the results in the previous section is that the output regulation problem cannot be directly reformulated as a classical PDE-ODE or ODE-PDE-ODE stabilization problem, where the (extended) system is mapped to a stable target system by adequately choosing the backstepping transform. This is due to the fact that the state of the system will not (in general) converge to the origin. There are some results using backstepping techniques that have a first component of motion planning and then use backstepping to track the desired trajectory, for instance [80]. In this section, we will focus instead on results based on output regulation and disturbance rejection techniques that do not require motion planning, since they are fundamentally different from stabilization problems (even around a reference trajectory).
@@
The first result in trajectory tracking based on a backstepping design can be found in [63], even if the main focus of the paper is the use of the adaptive techniques. In this paper, the system under consideration is an ODE with unknown input delay which, as has been previously discussed, is closely related to hyperbolic PDEs. A first result on disturbance rejection specifically for hyperbolic PDEs using the backstepping method can be found a few years later in [1], for linear hyperbolic systems including a boundary disturbance generated by an autonomous linear ODE (exosystem). More precisely, the boundary disturbance is considered as acting on the boundary opposite the control action, and the regulated output is a linear function of the PDE states on the same boundary as the disturbance (yet linearly independent). This result was later extended to a regulated output defined as a linear function of the PDE states at an arbitrary point inside the domain in [10], and subsequently to general heterodirectional transport equation systems with constant coefficients in [11]. The corresponding cases for spatially varying coefficients were studied in [109], [108] and [110]. In particular [108] links the corresponding design to a set of regulator equations depending on the localization of the regulated output. Two related results appearing in the same time period but on slightly different research lines are the disturbance rejection problem for 1D hyperbolic Partial Integro-Differential Equations (PIDEs) in [296] and the consideration of output regulation for anti-stable wave equations in [144] and coupled linear wave-ODE systems in [112]. Other recent results have focused a bit more on the trajectory tracking problem, such as [235] and [236] with a focus on linear hyperbolic systems and linear ODE exosystems, and [156] focusing on nonlinear ODEs.
@@
In a parallel track, some results have been obtained for analogous problems in the case of parabolic PDEs. In particular, [217] proposes a control design for the output-tracking problem of boundary actuated parabolic PDEs with varying parameters, using both a backstepping and flatness approach. In [107], the output regulation problem for a class of boundary-controlled parabolic PDEs is considered, where the disturbance can act both on the uncontrolled boundary of the PDE and the domain. Related to these results, other problems have also been considered, such as cooperative output-tracking problems that add the difficulty of time-varying structures, as in [112, 111]. It should be noted that several of the results focusing on output-tracking for parabolic systems via backstepping have followed developments in the hyperbolic cases, which probably explains the fact that papers in this area tend to be fewer and generally more consolidated results.
@@ 3.6 Coupled parabolic problems
Remarkably, the structure of these kernel equations is similar to the ones encountered in the backstepping control of coupled hyperbolic systems[155, 95]. This connection unveils a deeper underlying mathematical structure shared by backstepping designs for apparently different classes of PDE systems. Subsequently, [72] addressed the issue of boundary observer design for coupled reaction-diffusion systems with spatially varying reaction terms, utilizing a similar connection. Expanding on these advancements, [117] proposed a comprehensive backstepping design for coupled linear parabolic PIDEs (Partial Integro-Differential Equations) with distinct diffusion coefficients, spatially varying parameters, and mixed coupled boundary conditions. These work systematically tackled the well-posedness of the complex kernel equations, which constituted a system of coupled second-order hyperbolic PIDEs. By extending the method of successive approximations, the authors presented a constructive solution approach for the kernel equations, culminating in a robust control design methodology for a broader class of coupled parabolic systems, later expanded to time-varying coefficients in [166].
@@ 3.6.2 Couplings of parabolic PDEs with ODEs and chains
The interplay between parabolic PDEs and ODEs arises naturally in many applications, where actuator and sensor dynamics, or other lumped parameter subsystems, influence the behavior of systems governed by parabolic PDEs. Initial efforts, such as [177], considered the control of a single diffusion PDE coupled with a linear ODE. The authors employed a backstepping transformation incorporating both the PDE and ODE states, leading to a target system that could be stabilized using a boundary controller. The design focused on compensating the effect of the ODE dynamics on the PDE, achieving exponential stabilization of the combined system. A more complex scenario, involving a diffusion PDE sandwiched between two ODEs, was investigated in [258]. The system represented a plant with both actuator and sensor dynamics governed by ODEs, coupled with a diffusion process occurring within the plant. The control design involved a two-step backstepping procedure, addressing first the actuator dynamics and then the sensor dynamics. The resulting controller achieved stabilization of the sandwiched PDE and guaranteed the exponential convergence of the observer error.
Finally, [113] presented a comprehensive backstepping design for coupled parabolic ODE-PDE-ODE systems, where multiple parabolic PDEs with distinct diffusion coefficients and spatially varying parameters were bidirectionally coupled with ODEs at both boundaries. The authors employed a successive backstepping approach, systematically decoupling the system into a stable ODE-PDE-ODE cascade. The design involved solving a sequence of kernel equations, initial value problems, and boundary value problems, enabling the specification of the closed-loop stability margin. This work significantly broadened the scope of backstepping designs for interconnected parabolic-ODE systems, addressing a wider range of complex coupling structures and boundary conditions.
@@ 3.6.3 Hyperbolic-Parabolic Problems
The interconnection of hyperbolic and parabolic PDEs leads to mixed-class systems that capture the interplay between transport and diffusion phenomena. These systems arise in diverse applications, such as fluid-structure interactions, thermoacoustic instabilities, and chemotaxis models, where the finite speed of propagation inherent to hyperbolic dynamics interacts with the smoothing and dissipative effects of parabolic dynamics. Interestingly, the designs give rise to kernel equations that also exhibit a hyperbolic-parabolic structure. In particular, one of the backstepping kernel dealing with the parabolic state is defined on a square domain and verifies a parabolic equation, a unique feature not normally encountered in backstepping designs.
@@
A significant contribution towards tackling more general hyperbolic-parabolic PDE-PDE systems was made in [76]. This work addressed a system with boundary-dependent and in-domain direct coupling terms within the hyperbolic system domain, making controller design considerably more chal-lenging. The authors applied a novel backstepping transformation, leading to a set of coupled kernel equations that reflect the mixed hyperbolic-parabolic nature of the original system. The well-posedness of these coupled kernel equations was proven using an innovative infinite induction energy series, and the invertibility of the transformations was also established. This approach facilitated the design of a boundary controller that achieved exponential stabilization of the mixed PDE system in the sense.
@@ 3.6.4 Elliptic-parabolic problems
Elliptic-parabolic problems, arise in various applications where the steady-state behavior of one physical process (described by the elliptic equation) influences the transient behavior of another (described by the parabolic equation). While seemingly less common than purely parabolic or hyperbolic systems, these mixed-class systems present unique challenges for control design due to the inherent differences in their mathematical structure and underlying physical phenomena.
@@
The control of Navier-Stokes and magnetohydrodynamic (MHD) channel flow, studied respectively in [267] and [295], also present an elliptic-parabolic couplings. In both cases, the pressure verifies an elliptic PDE. In the case of MHD, there is an electrically conducting fluid with an imposed magnetic field which also verifies an elliptic PDE for the electric potential. Backstepping was applied to design a boundary controller that stabilized the velocity field by using the similar principle of solving the elliptic equation and using additional controls to make the solution strict-feedback and thus amenable to backstepping.
@@ 3.7.1 Parabolic case
The kernel satisfies a hyperbolic PDE in a hourglass-shaped domain instead of a triangle; this PDE can be split into two separate PDEs defined on triangular subdomains with coupled boundary conditions. The control laws are obtained from evaluating the transformation (73) at the boundaries, resulting in integral operators acting on the state :
@@
The bilateral backstepping design was also used in [53] to address the trajectory tracking problem for a class of viscous Hamilton-Jacobi PDEs, using bilateral actuation and sensing. The design employed a combination of feedback linearizing transformation and backstepping, resulting in explicit observer-based output-feedback laws that achieved local asymptotic stabilization of the desired trajectory. The feasibility conditions imposed by the linearizing transformation (its invertibility is only local) highlighted the regional nature of the stability result (see Section 4.4 for a similar unilateral result on the Burgers equation). A more recent paper [79] further explores the folding technique for designing bilateral backstepping output-feedback controllers for an unstable parabolic PDE. The authors introduce the notion of a folding point that can be arbitrarily chosen within the domain, leading to a more general formulation of the bilateral control problem. They also design a state observer based on collocated measurements at an arbitrary interior point, generating exponentially converging state estimates. The combined output feedback scheme, composed of the state-feedback and observer, is shown to stabilize the system. This paper also includes a numerical analysis of how the selection of folding point and measurement point affect the controller and observer responses, illustrating the potential for performance optimization using these design parameters.
@@ 3.8 Timoshenko beams and wave equations
Recently, designs for control of Timoshenko beams and wave equations using backstepping are relying on a preliminary step that involves a Riemann-like transformation to map the original system into a system of first-order hyperbolic PDEs, potentially coupled with ODEs. This transformation greatly facilitates the application of backstepping designs developed for interconnected systems that have been explored throughout this section, allowing for the design of controllers that achieve rapid stabilization, even in the presence of destabilizing boundary conditions and complex nonlocal terms.
@@ 4.1 Adaptive control and parameter estimation
As seen in the previous sections, the backstepping technique has provided constructive, and sometimes explicit, control laws for large classes of (mostly linear) PDE systems. It has also enabled the use of other advanced methods and multi-step design techniques by simplifying the structure of closed-loop PDE systems. One of the first such uses of the backstepping methodology was the extension of adaptive control designs to new classes of PDEs. These designs leveraged the simpler structure of target systems obtained via the backstepping transform and, when available, the explicit nature of the transform, in order to apply different adaptive strategies to the control of both parabolic and hyperbolic PDEs. These strategies also allow for the estimation of boundary parameters in certain classes of systems and are not limited to in-domain parameters. The use of backstepping-based adaptive designs appeared in [249], applied to 1D parabolic PDEs and was rapidly expanded in the series of papers [191, 250, 251], as well as extended to higher-dimensional parabolic PDEs in [192]. These and other designs were collected in the book [253]. In the case of delay systems, closely related to hyperbolic PDE systems, the first results for unknown delay were obtained in [183, 64] for and ODE with unknown input delay and extended to the problem of trajectory tracking for an output of the ODE in [63]. As for hyperbolic PDEs, the first result is [180], followed a few years later by [55] for a 1D transport PDE with non-local terms. Closely thereafter, an adaptive output-feedback control law for a wave equation with dynamic boundary appeared in [65]. After these initial results, many papers on adaptive schemes based on the backstepping technique have appeared in the literature, for instance, for general linear hyperbolic PDEs in [16]. A good overview of these designs can be found in [13].
@@ 4.2 Extremum seeking for PDEs
In the book [223], Oliveira and Krstic assemble a multitude of algorithms for seeking extrema and Nash equilibria (in non-cooperative multi-agent scenarios) in the presence of parabolic and hyperbolic PDEs, including delays. While the presence of stable PDEs can be ignored by resorting to sufficiently slow seeking, when the PDEs are compensated with PDE backstepping the limit on the convergence speed is removed. And the convergence rate is assignable with Newton algorithms.
@@ 4.3 Extension to higher-dimensional domains
For many models in fluid dynamics and other areas, the design of boundary controllers and observers can be greatly simplified by exploiting the concept of spatial invariance [42]. This property, also known as translational invariance, arises when the system dynamics and geometry are invariant under translations in one or more spatial coordinates. This allows for a reduction in the system’s dimensionality by transforming the spatially invariant coordinates into a parameter or set of parameters. The controllers and observers can then be designed for these simpler systems and subsequently reconstructed in the original physical space. This procedure requires that the system possess specific properties, namely that the spatially invariant coordinates lie within a group (usually the real line or a periodic domain), actuation and sensing are fully distributed over those coordinates, and the dynamics and geometry are invariant with respect to translations in them. By applying tools like the Fourier transform (for infinite domains) or Fourier series/spherical harmonics (for periodic domains), one can convert the spatially invariant coordinates into wave numbers or harmonics, transforming spatial derivatives into algebraic multiplications. This simplifies the control design and analysis, allowing for the application of techniques like backstepping to each wave number separately. The earliest designs based on these ideas were for flow control applications, and are explored in Section 5.3. Later, [274] and [276] introduced backstepping designs for reaction-diffusion equations on, respectively, a 2-D disk and a 3-D sphere. The authors exploited the radial symmetry of these domains to simplify the kernel equations and obtained explicit control laws that achieved exponential stabilization. These results were generalized in [281] to address the control of reaction-diffusion equations on a n-dimensional ball, employing spherical harmonics and Bessel functions to derive the backstepping kernel and feedback gains. Interestingly, the 1-dimensional case can be exploited to obtain the bilateral design described in Section 3.7.1. More recently, [280] presented a backstepping design for a reaction-diffusion system in a ball of arbitrary dimension with radially varying reaction using the power series method to overcome the singularities in the kernel equations. Other works exploring higher-dimensional designs include [208] which exploits simmetry to apply the classic backstepping designs to simple geometries, and the book [215], which provides a comprehensive overview of backstepping designs for higher-dimensional PDEs in several domains, including those for plates, shells, and other flexible structures.

### v3

@@ 3.4 Interconnected systems with a chain structure
The framework proposed by the representation (51) is comprehensive and can describe various interconnected systems. As shown in [21], using a technique referred to as folding (see [104, 20, 25] for details), it can describe complex networks composed of interconnected hyperbolic/ODEs subsystems. However, such a representation shadows the cascade structure between the different subsystems, which may lead to strong requirements when designing stabilizing controllers. In particular, the contributions presented in Section 3.3.3 avoid configurations where the PDE system is under-actuated or under-measured. Consequently, they fail to design stabilizing controllers for interconnected PDE systems when only one of the PDEs is actuated. Consider the following interconnections of PDE of heterodirectional systems
@@ 3.5 Output regulation and disturbance rejection
A natural extension of the analysis of coupled PDE-ODE systems is the consideration of output tracking and disturbance rejection problems with a disturbance or reference model. In this case, when the disturbance and output trajectories can be modeled by a finite-dimensional exosystem, backstepping designs exist for the disturbance rejection and output tracking problems. The main technical difference with respect to the results in the previous section is that the output regulation problem cannot be directly reformulated as a classical PDE-ODE or ODE-PDE-ODE stabilization problem, where the (extended) system is mapped to a stable target system by adequately choosing the backstepping transform. This is due to the fact that the state of the system will not (in general) converge to the origin. There are some results using backstepping techniques that have a first component of motion planning and then use backstepping to track the desired trajectory, for instance, [81]. In this section, we will focus instead on results based on output regulation and disturbance rejection techniques that do not require motion planning, since they are fundamentally different from stabilization problems (even around a reference trajectory).
@@
The first result in trajectory tracking based on a backstepping design can be found in [64], even if the main focus of the paper is the use of adaptive techniques. In this paper, the system under consideration is an ODE with unknown input delay which, as has been previously discussed, is closely related to hyperbolic PDEs. A first result on disturbance rejection specifically for hyperbolic PDEs using the backstepping method can be found a few years later in [1], for linear hyperbolic systems including a boundary disturbance generated by an autonomous linear ODE (exosystem). More precisely, the boundary disturbance is considered as acting on the boundary opposite the control action, and the regulated output is a linear function of the PDE states on the same boundary as the disturbance (yet linearly independent). This result was later extended to a regulated output defined as a linear function of the PDE states at an arbitrary point inside the domain in [10], and subsequently to general heterodirectional transport equation systems with constant coefficients in [11]. The corresponding cases for spatially varying coefficients were studied in [110], [109] and [111]. In particular [109] links the corresponding design to a set of regulator equations depending on the localization of the regulated output. Two related results appearing in the same time period but on slightly different research lines are the disturbance rejection problem for 1D hyperbolic Partial Integro-Differential Equations (PIDEs) in [300] and the consideration of output regulation for anti-stable wave equations in [145] and coupled linear wave-ODE systems in [113]. Other recent results have focused a bit more on the trajectory tracking problem, such as [236] and [237] with a focus on linear hyperbolic systems and linear ODE exosystems, and [157] focusing on nonlinear ODEs.
@@
In a parallel track, some results have been obtained for analogous problems in the case of parabolic PDEs. In particular, [218] proposes a control design for the output-tracking problem of boundary-actuated parabolic PDEs with varying parameters, using both a backstepping and flatness approach. In [108], the output regulation problem for a class of boundary-controlled parabolic PDEs is considered, where the disturbance can act both on the uncontrolled boundary of the PDE and the domain. Related to these results, other problems have also been considered, such as cooperative output-tracking problems that add the difficulty of time-varying structures, as in [113, 112]. It should be noted that several of the results focusing on output-tracking for parabolic systems via backstepping have followed developments in the hyperbolic cases, which probably explains the fact that papers in this area tend to have fewer and generally more consolidated results.
@@ 3.6 Coupled parabolic problems
Remarkably, the structure of these kernel equations is similar to the ones encountered in the backstepping control of coupled hyperbolic systems[156, 96]. This connection unveils a deeper underlying mathematical structure shared by backstepping designs for apparently different classes of PDE systems. Subsequently, [73] addressed the issue of boundary observer design for coupled reaction-diffusion systems with spatially varying reaction terms, utilizing a similar connection. Expanding on these advancements, [118] proposed a comprehensive backstepping design for coupled linear parabolic PIDEs (Partial Integro-Differential Equations) with distinct diffusion coefficients, spatially varying parameters, and mixed coupled boundary conditions. This work systematically tackled the well-posedness of the complex kernel equations, which constituted a system of coupled second-order hyperbolic PIDEs. By extending the method of successive approximations, the authors presented a constructive solution approach for the kernel equations, culminating in a robust control design methodology for a broader class of coupled parabolic systems, later expanded to time-varying coefficients in [167].
@@ 3.6.2 Couplings of parabolic PDEs with ODEs and chains
The interplay between parabolic PDEs and ODEs arises naturally in many applications, where actuator and sensor dynamics, or other lumped parameter subsystems, influence the behavior of systems governed by parabolic PDEs. Initial efforts, such as [178], considered the control of a single diffusion PDE coupled with a linear ODE. The authors employed a backstepping transformation incorporating both the PDE and ODE states, leading to a target system that could be stabilized using a boundary controller. The design focused on compensating the effect of the ODE dynamics on the PDE, achieving exponential stabilization of the combined system. A more complex scenario involving a diffusion PDE sandwiched between two ODEs was investigated in [260]. The system represented a plant with both actuator and sensor dynamics governed by ODEs, coupled with a diffusion process occurring within the plant. The control design involved a two-step backstepping procedure, addressing first the actuator dynamics and then the sensor dynamics. The resulting controller achieved stabilization of the sandwiched PDE and guaranteed the exponential convergence of the observer error.
Finally, [114] presented a comprehensive backstepping design for coupled parabolic ODE-PDE-ODE systems, where multiple parabolic PDEs with distinct diffusion coefficients and spatially varying parameters were bidirectionally coupled with ODEs at both boundaries. The authors employed a successive backstepping approach, systematically decoupling the system into a stable ODE-PDE-ODE cascade. The design involved solving a sequence of kernel equations, initial value problems, and boundary value problems, enabling the specification of the closed-loop stability margin. This work significantly broadened the scope of backstepping designs for interconnected parabolic PDEs and ODE systems, addressing a wider range of complex coupling structures and boundary conditions.
@@ 3.6.3 Hyperbolic-Parabolic Problems
The interconnection of hyperbolic and parabolic PDEs leads to mixed-class systems that capture the interplay between transport and diffusion phenomena. These systems arise in diverse applications, such as fluid-structure interactions, thermoacoustic instabilities, and chemotaxis models, where the finite speed of propagation inherent to hyperbolic dynamics interacts with the smoothing and dissipative effects of parabolic dynamics. Interestingly, the designs give rise to kernel equations that also exhibit a hyperbolic-parabolic structure. In particular, one of the backstepping kernels dealing with the parabolic state is defined on a square domain and verifies a parabolic equation, a unique feature not normally encountered in backstepping designs.
@@
A significant contribution towards tackling more general hyperbolic-parabolic PDE-PDE systems was made in [77]. This work addressed a system with boundary-dependent and in-domain direct coupling terms within the hyperbolic system domain, making controller design considerably more challenging. The authors applied a novel backstepping transformation, leading to a set of coupled kernel equations that reflect the mixed hyperbolic-parabolic nature of the original system. The well-posedness of these coupled kernel equations was proven using an innovative infinite induction energy series, and the invertibility of the transformations was also established. This approach facilitated the design of a boundary controller that achieved exponential stabilization of the mixed PDE system in the sense.
@@ 3.6.4 Elliptic-parabolic problems
Elliptic-parabolic problems arise in various applications where the steady-state behavior of one physical process (described by the elliptic equation) influences the transient behavior of another (described by the parabolic equation). While seemingly less common than purely parabolic or hyperbolic systems, these mixed-class systems present unique challenges for control design due to the inherent differences in their mathematical structure and underlying physical phenomena.
@@
The control of Navier-Stokes and magnetohydrodynamic (MHD) channel flows, studied respectively in [270] and [299], also present elliptic-parabolic couplings. In both cases, the pressure verifies an elliptic PDE. In the case of MHD, there is an electrically conducting fluid with an imposed magnetic field which also verifies an elliptic PDE for the electric potential. Backstepping was applied to design a boundary controller that stabilized the velocity field by using the similar principle of solving the elliptic equation and using additional controls to make the solution strict-feedback and thus amenable to backstepping.
@@ 3.7.1 Parabolic case
The kernel satisfies a hyperbolic PDE in an hourglass-shaped domain instead of a triangle; this PDE can be split into two separate PDEs defined on triangular subdomains with coupled boundary conditions. The control laws are obtained from evaluating the transformation (73) at the boundaries, resulting in integral operators acting on the state :
@@
The bilateral backstepping design was also used in [53] to address the trajectory tracking problem for a class of viscous Hamilton-Jacobi PDEs, using bilateral actuation and sensing. The design employed a combination of feedback linearizing transformation and backstepping, resulting in explicit observer-based output-feedback laws that achieved local asymptotic stabilization of the desired trajectory. The feasibility conditions imposed by the linearizing transformation (its invertibility is only local) highlighted the regional nature of the stability result (see Section 4.4 for a similar unilateral result on the Burgers equation). A more recent paper [80] further explores the folding technique for designing bilateral backstepping output-feedback controllers for an unstable parabolic PDE. The authors introduce the notion of a folding point that can be arbitrarily chosen within the domain, leading to a more general formulation of the bilateral control problem. They also design a state observer based on collocated measurements at an arbitrary interior point, generating exponentially converging state estimates. The combined output feedback scheme, composed of the state-feedback and observer, is shown to stabilize the system. This paper also includes a numerical analysis of how the selection of folding points and measurement points affect the controller and observer responses, illustrating the potential for performance optimization using these design parameters.
@@ 3.8 Timoshenko beams and wave equations
Recent designs for control of Timoshenko beams and wave equations using backstepping rely on a preliminary step that involves a Riemann-like transformation to map the original system into a system of first-order hyperbolic PDEs, potentially coupled with ODEs. This transformation greatly facilitates the application of backstepping designs developed for interconnected systems that have been explored throughout this section, allowing for the design of controllers that achieve rapid stabilization, even in the presence of destabilizing boundary conditions and complex nonlocal terms.
@@ 4.1 Adaptive control and parameter estimation
As seen in the previous sections, the backstepping technique has provided constructive and sometimes explicit control laws for large classes of (mostly linear) PDE systems. It has also enabled the use of other advanced methods and multi-step design techniques by simplifying the structure of closed-loop PDE systems. One of the first such uses of the backstepping methodology was the extension of adaptive control designs to new classes of PDEs. These designs leveraged the simpler structure of target systems obtained via the backstepping transform and, when available, the explicit nature of the transform, in order to apply different adaptive strategies to the control of both parabolic and hyperbolic PDEs. These strategies also allow for the estimation of boundary parameters in certain classes of systems and are not limited to in-domain parameters. The use of backstepping-based adaptive designs appeared in [251], applied to 1D parabolic PDEs and was rapidly expanded in the series of papers [192, 252, 253], as well as extended to higher-dimensional parabolic PDEs in [193]. These and other designs were collected in the book [255]. In the case of delay systems, closely related to hyperbolic PDE systems, the first results for unknown delay were obtained in [184, 65] for and ODE with unknown input delay and extended to the problem of trajectory tracking for an output of the ODE in [64]. As for hyperbolic PDEs, the first result is [181], followed a few years later by [55] for a 1D transport PDE with non-local terms. Closely thereafter, an adaptive output-feedback control law for a wave equation with dynamic boundary appeared in [66]. After these initial results, many papers on adaptive schemes based on the backstepping technique have appeared in the literature, for instance, for general linear hyperbolic PDEs in [16]. A good overview of these designs can be found in [13].
@@ 4.2 Extremum seeking for PDEs
In the book [224], Oliveira and Krstic assemble a multitude of algorithms for seeking extrema and Nash equilibria (in non-cooperative multi-agent scenarios) in the presence of parabolic and hyperbolic PDEs, including delays. While the presence of stable PDEs can be ignored by resorting to sufficiently slow seeking, when the PDEs are compensated with PDE backstepping the limit on the convergence speed is removed and the convergence rate is assignable with Newton algorithms.
@@ 4.3 Extension to higher-dimensional domains
For many models in fluid dynamics and other areas, the design of boundary controllers and observers can be greatly simplified by exploiting the concept of spatial invariance [42]. This property, also known as translational invariance, arises when the system dynamics and geometry are invariant under translations in one or more spatial coordinates. This allows for a reduction in the system’s dimensionality by transforming the spatially invariant coordinates into a parameter or set of parameters. The controllers and observers can then be designed for these simpler systems and subsequently reconstructed in the original physical space. This procedure requires that the system possess specific properties, namely that the spatially invariant coordinates lie within a group (usually the real line or a periodic domain), actuation and sensing are fully distributed over those coordinates, and the dynamics and geometry are invariant with respect to translations in them. By applying tools like the Fourier transform (for infinite domains) or Fourier series/spherical harmonics (for periodic domains), one can convert the spatially invariant coordinates into wave numbers or harmonics, transforming spatial derivatives into algebraic multiplications. This simplifies the control design and analysis, allowing for the application of techniques like backstepping to each wave number separately. The earliest designs based on these ideas were for flow control applications, and are explored in Section 5.3. Later, [277] and [280] introduced backstepping designs for reaction-diffusion equations on, respectively, a 2-D disk and a 3-D sphere. The authors exploited the radial symmetry of these domains to simplify the kernel equations and obtained explicit control laws that achieved exponential stabilization. These results were generalized in [278] to address the control of reaction-diffusion equations on a n-dimensional ball, employing spherical harmonics and Bessel functions to derive the backstepping kernel and feedback gains. Interestingly, the 1-dimensional case can be exploited to obtain the bilateral design described in Section 3.7.1. More recently, [285] presented a backstepping design for a reaction-diffusion system in a ball of arbitrary dimension with radially varying reaction using the power series method to overcome the singularities in the kernel equations. Other works exploring higher-dimensional designs include [209], which exploits symmetry to apply the classic backstepping designs to simple geometries, and the book [216], which provides a comprehensive overview of backstepping designs for higher-dimensional PDEs in several domains, including those for plates, shells, and other flexible structures.